OpenStudy (anonymous):

Calculus A cylinder shaped can needs to be constructed to hold 350 cubic centimeters of soup. The material for the sides of the can costs 0.02 cents per square centimeter. The material for the top and bottom of the can need to be thicker, and costs 0.05 cents per square centimeter. Find the dimensions for the can that will minimize production cost. Volume of a cylinder: V = pi r^2 h Area of the sides: A = 2pi r h Area of the top/bottom: A = pi r^2 To minimize the cost of the can: Radius of the can: Height of the can: Minimum Cost

OpenStudy (anonymous):

Lets solve for h first. \[V= \pi r^2h \implies 350 cm^3 = \pi r^2 h \implies h = \frac{ 350cm^3 }{ \pi r^2 } \]

OpenStudy (anonymous):

Following

OpenStudy (anonymous):

Now we can also find the surface area

OpenStudy (anonymous):

by doing the cost function?

OpenStudy (anonymous):

\[S = 2 \pi r h+2 \pi r^2\] so solve for this

OpenStudy (anonymous):

I did not know this part

OpenStudy (anonymous):

You'll have to plug in for h in the second equation and then you can find cost

OpenStudy (anonymous):

S=2πr(350cm3πr2)+2πr

OpenStudy (anonymous):

Idk how to do the Pi smybol i tried copy and paste

OpenStudy (anonymous):

Yeah I can't see it, only shows question marks.

OpenStudy (anonymous):

so sub the H with the height equation?

OpenStudy (anonymous):

Yes or you can just find the integer h but, \[S = 2 \pi r (\frac{ 350 cm^3 }{ \pi r ^2 })+2 \pi r^2\]

OpenStudy (anonymous):

ok which is what I got.

OpenStudy (anonymous):

Yeah just simplify it and what do you get?

OpenStudy (anonymous):

Then we can find the cost of the materials

OpenStudy (anonymous):

2(pi)r*350

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

\[S = 2\frac{ 350 cm^3 }{ r } + 2 \pi r^2\]

OpenStudy (anonymous):

I subtracted the 2(pi)r and then multiplied by (pi)r^2 to get 2(pi)r^3

OpenStudy (anonymous):

Now the cost\[C = 0.02\left( \frac{ 700 }{ r } \right) + 0.05\left( 2 \pi r^2 \right)\] Solve for this and find the derivative

OpenStudy (anonymous):

Once you find the derivative set it to 0 and solve for r. So you'll have h and r then go back and plug all the numbers in for h and you'll have the minimum cost.

OpenStudy (anonymous):

Oh ok, that makes sense now, so what did I do wrong with the equation to get to where you got?

OpenStudy (anonymous):

I think you just plugged in the wrong numbers, it's hard to tell as it's not in LaTeX, so I really didn't read your equation haha. If you drew it out, I can see what you did wrong :d

OpenStudy (anonymous):

Oh I see, you were multiplying when you should be dividing.

OpenStudy (anonymous):

Let me give it a go

OpenStudy (anonymous):

Take your time

OpenStudy (anonymous):

\[2\pi(r)(h)+2\pi(r)=2\pi(r)(350(\div)\pi(r)^2)-2\pi(r)^2\]

OpenStudy (anonymous):

So here I subtracted the 2pi(r) to bring over to the left, then multiplied by pi(r)^2 to get 2pi(r)^3 on the left and 2pi(r)2 on the left

OpenStudy (anonymous):

Why are you doing that?

OpenStudy (anonymous):

Those are two separate equations you don't have to do that, I was just showing solve for one equation (finding h) then plug that into the surface area equation.

OpenStudy (anonymous):

You over complicated the problem :), no worries though, do you understand now?

OpenStudy (anonymous):

Oh Ok yeah I understand now, the benefits of calculus is me always doing that haha but yes I now understand thank you very much!

OpenStudy (anonymous):

Np :)